# Toward a unified theory for common affine roots of general sets of multivariate polynomials

### v2

@@ Abstract
For univariate polynomials over arbitrary field the degree gives an upper bound on the number of roots (factor theorem) and as a related result for any finite point-set one can construct a polynomial of degree equal to the cardinality having all the points as roots (interpolation theorem). Terence Tao noted in [39] that the theory of multivariate polynomials is not yet sufficiently matured to provide similar theorems with an equally simple relation between them. In the present paper we argue that for general multivariate polynomials the right measure for the size of the polynomial should not be the degree, but the leading monomial. In this setting the footprint bound [23] becomes a natural enhancement of the factor theorem providing a bound on the number of common roots of general multivariate polynomials which is sharp for all finite Cartesian product point sets. Furthermore, by using methods from the theory of error-correcting codes we establish a natural formulation of the interpolation theorem to the case of common roots of multivariate polynomials. In short the two theorems reduce to the same result, but for dual spaces, establishing the unification requested in [39]. We leave it for further research to possibly establish similar related results taking multiplicity into account.
Keywords: Affine roots, footprint bound, generalized Alon-Füredi bound, interpolation, multivariate polynomials
@@ 1 Introduction
For a univariate non-zero polynomial over an arbitrary field the number of roots is at most equal to its degree, and if the field is algebraically closed and if roots are counted with multiplicity then equality holds. This result is in close correspondence with the fact that the interpolation problem of finding a polynomial of lowest possible degree evaluating to 0 on a finite point-set boils down to calculating the product of linear terms each having one of the desired elements as root, and by repeating terms one can also deal with multiplicity. The situation is very much in contrast to that of multivariate polynomials. Firstly, a multivariate polynomial ring is not a unique factorization domain, and therefore the considered interpolation problem is not about multiplying linear factors. Secondly, a non-constant multivariate polynomial over an infinite field has infinitely many roots, and therefore if the aim is to consider a case of finitely many roots then one either needs to restrict to roots from a prescribed finite point-set or one must study mutual roots of a set of polynomials which together define a zero-dimensional ideal.
@@
    Given some measure on the sizes of polynomials let $a,k$ be any positive integers. If we are to guarantee that for any point-set $A$ of cardinality at most $a$ there exist $k$ polynomials with pairwise different leading monomials having $A$ as common roots, what is then the sharpest upper bound in terms of the measure on sizes of polynomials we can formulate?
Our first step in this direction is to replace in our analysis the concept of degree with that of leading monomial being then the chosen measure in the present paper. For univariate polynomials there is only one choice of monomial ordering and the leading monomial of a polynomial $F(X)$ of degree $d\geq 0$ equals $X^{d}$, hence the two concepts are basically the same. In contrast, for multivariate polynomials there are infinitely many monomial orderings and for many of these the total degree of the leading monomial does not even need to be equal to the total degree of the polynomial under consideration. Considering leading monomial(s) is not only as natural a choice for our study as considering total degree(s); as shall be demonstrated it is the key to achieve the above mentioned enhancement and connection. Even more importantly, using the leading monomial rather than the total degree provides much more fine-grained information which is sharp in more cases.
In his exposition [39] Terence Tao treats problems (I) and (II) from the above paragraph including discussing on the one side Alon’s famous combinatorial Nullstellensatz [1], the Schwartz-Zippel bound [42, 36, 29] (as well as Dvir’s generalization taking multiplicity into account [13, 14, 15]) and on the other side the below standard result [39][Lem. 1] on interpolation; with the aim of solving problems in extremal combinatorics:
@@
Let $\mathbb{F}$} be a field, let $m\geq 1$ be an integer, and $d\geq 0$. If $A\subset{\mathbb{F}}^{m}$ has cardinality less than ${d+m}\choose{m}$, then there exists a non-zero polynomial $F\in{\mathbb{F}}[X_{1},\ldots,X_{n}]$ of degree at most $d$ having $A$ as roots.
@@
The theory of commutative algebra and algebraic geometry is very rich and indeed Tao in addition to the above results treats a cornucopia of deep results such as the Alon-Füredi bound, Bezout’s Theorem, Stepanov’s Method, the Hasse Bound and many more [4, 8, 27, 40, 30, 37, 41], as well as provide interpolation counterparts in more cases. Our contribution is to address the problem of common roots of general polynomials where the only assumption is that no two of them have the same leading monomial which is no restriction as we can always apply the division algorithm [10][Sec, 2, Par. 3]] without effecting the set of common roots. Similarly, we consider any finite point-sets without any restriction on their structure. Therefore, our answer to (II) should be compared to Proposition 1 and our answer to (I) should be compared to the Schwartz-Zippel bound, the Alon-Füredi bound and the generalized Alon-Füredi bound [7]. In both cases the material of the present paper can be viewed as improvements and of course also as enhancements as we do not restrict to single polynomials but rather consider arbitrary finite set of polynomials.
@@
In the present paper we give refined answers to both (I) and (II) taking the leading monomial into account and we enhance the methods to treat the case of common roots of a set of more polynomials. Knowing the leading monomial corresponds to having important information regarding the most crucial coefficient and thereby our insight addresses the remark by Tao. More importantly, we demonstrate a very strong relationship between our answers to (I) and (II) which is then our proposal for the requested useful unification, except we do not treat multiplicity. We leave it for further research to explore implications in the theory of extremal combinatorics.
At the heart of our method is insight from the theory of error-correcting codes which we translate to hold for general fields rather than just for finite fields. Our answer to (I) is basically to recall the footprintbound [23] in some of its many variants including enhancing a formulation known as the Feng-Rao bound for primary codes [3, 20, 24] to hold for arbitrary field. When applied to finite Cartesian product point-sets the footprint bound is sharp in that it is easy to find a polynomial which is the product of linear terms having exactly as many roots as predicted by the bound. Our answer to (II) is almost trivial in the case of only one monomial, but to treat common roots of more polynomials we shall need some basic lemmas by Forney [19] on linear codes over arbitrary field. The answer to (II) is then about applying an enhanced version of the Feng-Rao bound for dual codes [16, 17, 18, 28, 25], which then also demonstrates the relationship between the two methods for arbitrary fields as well as demonstrate the sharpness of our answer to (II) in the case of finite fields.
@@
The paper is organized as follows. We start in Section 2 by presenting our main finding regarding (II) and by recalling the footprint bound which addresses (I). In that section we also provide a first description of the relationship between these results. To prepare for the proof of the first mentioned result we then in Section 3 treat related problems in the language of error-correcting codes over arbitrary fields. In Section 4 as a corollary we then obtain the proof of our answer to (II) which also demonstrates the strong relationship between (I) and (II), and which implies that our answer to (II) is sharp in the case of finite fields. Section 5 contains concluding remarks.
@@ 2 Main results on multivariate polynomials
Our main result is the following enhancement of Proposition 1 which surprisingly is best possible as for finite fields we can show it to be sharp. Observe that if $\mathbb{F}$} is a finite field with $q$ elements, in which case we write ${\mathbb{F}}={\mathbb{F}}_{q}$, and if $q\leq d$ then Proposition 1 does not give us any real information as all points in ${\mathbb{F}}^{n}$ are then trivially roots of each of the polynomials $X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}$. The below theorem is prepared to also give non-trivial insight in such cases.
@@
For multivariate monomials $N,N_{1},\ldots,N_{s}$ write
$\mu(N)$ $:=$ $\#\{M\mid M{\mbox{ is a monomial which divides }}N\}$
@@
constitutes a basis for ${\mathbb{F}}_{q}^{n}$ as a vector-space over ${\mathbb{F}}_{q}$. In particular for any polynomial $G$ with support not being a subset of $\Box$ there exists a polynomial $H$ with support being a subset of $\Box$ in such a way that $G$ and $H$ evaluates similarly at ${\mathbb{F}}_{q}^{m}$ and such that ${\mbox{lm}}(H)\preceq{\mbox{lm}}(G)$ holds for any choice of monomial ordering $\prec$. Here, ${\mbox{lm}}()$ denotes the leading monomial. More concretely, $H$ can be obtained from $G$ by dividing the latter with $\{X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}\}$. Therefore it is no restricting to consider in the finite field case only polynomials with support in $\Box$.
@@
We note that the right hand side of (2) equals the dimension of the q-ary Reed-Muller code ${\mathcal{RM}}_{q}(d,m)$ [12, Lem. 2.4.6]).
We next address (I) by recalling the footprint bound on the number of common roots of any given set of polynomials over arbitrary field [23]. To the best of our knowledge this is the most general and sharp bound for arbitrary sets of polynomials over arbitrary point-sets. Let $\prec$ be a monomial ordering on the set of monomials in $m$ variables, and consider an ideal $I\subseteq{\mathbb{F}}[X_{1},\ldots,X_{m}]$. The footprint of $I$ with respect to $\prec$ is
@@
The crucial observation is that $\{M+I\mid M\in\Delta_{\prec}(I)\}$ constitutes a basis for ${\mathbb{F}}[X_{1},\ldots,X_{m}]/I$ as a vector-space over $\mathbb{F}$} which is seen by applying the division algorithm.
@@
Assume first that $V_{\mathbb{F}}(I)$ is finite and write $V_{\mathbb{F}}(I)=\{\beta_{1},\ldots,\beta_{n}\}$. Consider the evaluation map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(\beta_{1}),\ldots,F(\beta_{n}))$. Due to Hermite interpolation this map is surjective and the first part of the theorem follows. To see the last part we use the fact [10][Ch. 4, Sec. 2, Th. 7] that ${\mathcal{I}}(V_{\mathbb{F}}(I))=I$ whenever $I$ is radical and $\mathbb{F}$} is algebraically closed. Aiming for a contradiction we finally assume that $V_{\mathbb{F}}(I)$ is infinite and choose $\{\beta_{1},\ldots,\beta_{n}\}\subseteq V_{\mathbb{F}}(I)$ with $n>\#\Delta_{\prec}(I)$. But then ev cannot be surjective. ∎
@@
$I=\left\langle G_{1}=\prod_{b\in A_{1}}(X_{1}-b),\ldots,G_{m}=\prod_{b\in A_{m}}(X_{m}-b)\right\rangle$
@@
$\Delta_{\prec}(I)=\{M{\mbox{ is a monomial }}\mid\deg_{X_{1}}M<\#A_{1},\ldots,\deg_{X_{m}}M<\#A_{m}\}=:\Diamond.$
Note, that if $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$, then $\Diamond=\Box$ where the latter is described in Definition 3.
@@
Let $A_{1}\times\cdots\times A_{m}$ be a finite point-set and consider pairwise different monomials $M_{1},\ldots,M_{t}$ in $\Diamond$. Let $\prec$ be an arbitrary monomial ordering. For general set of $k\leq t$ polynomials $F_{1},\ldots,F_{k}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with ${\mbox{Supp}}(F_{i})\in\Diamond$, $i=1,\ldots,k$ and leading monomials being pairwise different all belonging to $\{M_{1},\ldots,M_{t}\}$ the number of non-common roots of $F_{1},\ldots,F_{k}$ in $A_{1}\times\cdots\times A_{m}$ is at least
@@
where $N=X_{1}^{\#A_{1}-1}\cdots X_{m}^{\#A_{m}-1}$. This bound is sharp.
@@
We only need to demonstrate sharpness. To this end write $a_{i}=\#A_{i}$, $A_{i}=\{b_{1}^{(i)},\ldots,b_{a_{i}}^{(i)}\}$ for $i=1,\ldots,m$. To each monomial $N\in\Diamond$ we associate the polynomial
@@
which clearly has leading monomial $N$. Then for any set of monomials $N_{1},\ldots,N_{s}\subseteq\Diamond$ the number of non-common roots of $H_{N_{1}},\ldots,H_{N_{s}}$ is precisely $\sigma(N_{1},\ldots,N_{s})$. ∎
A special case of Corollary 10 is when $A_{1}=\cdots=\#A_{m}=:a$ and there is only one polynomial $F$, the only information of which we have is its total degree $\deg F=d$. Writing $d=(\sum_{i=1}^{v}(a-1))+\ell$ where $0\leq\ell<a-1$ we have
@@
which is a special case of the Alon-Füredi bound [2] which follows from Corollary 10 in a similar manner.
@@
Let ${\mathbb{F}}={\mathbb{F}}_{8}$, i.e. the field with $8$ elements, and consider $\langle X_{1}^{8}-X_{1},X_{2}^{8}-X_{2}\rangle\subseteq{\mathbb{F}}_{8}[X_{1},X_{2}]$. Clearly, $V_{\mathbb{F}_{8}}(I)={\mathbb{F}}_{8}^{2}$. We have $\Delta_{\prec}(I)=\{X_{1}^{i_{1}}X_{2}^{i_{2}}\mid 0\leq i_{1},i_{2}<8\}$ regardless of the choice of monomial ordering $\prec$. This set of course is the set $\Box$ from Definition 3. Assume $M_{1}=X_{1}^{3}X_{2}$, $M_{2}=X_{1}^{2}X_{2}^{2}$, and $M_{3}=X_{1}X_{2}^{3}$ and that $\prec$ is the degree lexicographic ordering with $X_{1}\prec X_{2}$ which implies that indeed $M_{1},M_{2},M_{3}$ are strictly consecutive. We have $\mu(M_{1})=\mu(M_{3})=8$, $\mu(M_{2})=9$, $\mu(M_{1},M_{2})=\mu(M_{2},M_{3})=11$, $\mu(M_{1},M_{3})=12$, and $\mu(M_{1},M_{2},M_{3})=13$. From this we conclude that for each $A\subseteq{\mathbb{F}}_{8}^{2}$ of size at most $7$ there exist polynomials $F_{1},F_{2},F_{3}$, ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, ${\mbox{lm}}(F_{3})=M_{3}$ having all elements of $A$ as common roots. We similarly conclude that whenever $A$ is of size at most $10$ then there exist $F_{1},F_{2}$ with ${\mbox{lm}}(F_{1})\neq{\mbox{lm}}(F_{2})$, ${\mbox{lm}}(F_{1}),{\mbox{lm}}(F_{2})\in\{M_{1},M_{2},M_{3}\}$ with all elements of $A$ as common roots. Finally for any $A$ with #A<13\#A<13 there exists $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{2},M_{3}\}$ such that $F(a)=0$ for all $a\in A$. In the other direction $\sigma(M_{1})=\sigma(M_{3})=35$, $\sigma(M_{2})=36$, $\sigma(M_{1},M_{2})=\sigma(M_{2},M_{3})=42$, $\sigma(M_{1},M_{3})=45$ and $\sigma(M_{1},M_{2},M_{3})=46$ which tells us that any polynomial $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{3}\}$ has at most 6435=2964-35=29 roots and that if ${\mbox{lm}}(F)=M_{2}$ then it has at most 6436=2864-36=28 roots. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{2}$ or if ${\mbox{lm}}(F_{1})=M_{2}$ and ${\mbox{lm}}(F_{2})=M_{3}$ then $F_{1}$ and $F_{2}$ have at most 6442=2264-42=22. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{3}$ the number of common roots is at most 6445=1964-45=19. Finally $F_{1},F_{2},F_{3}$ with ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, and ${\mbox{lm}}(F_{3})=M_{3}$ have at most 6446=1864-46=18 roots in common.
The similarity in expression (1) and expression (3) suggests that there is a strong relationship between Theorem 5 and Corollary 10 which we now explain.

### v3

@@ Abstract
For univariate polynomials over arbitrary field the degree gives an upper bound on the number of roots (factor theorem) and as a related result for any finite point-set one can construct a polynomial of degree equal to the cardinality having all the points as roots (interpolation theorem). Tao noted in [44] that the theory of multivariate polynomials is not yet sufficiently matured to provide similar theorems with an equally simple relation between them. In the present paper we argue that for general multivariate polynomials the right measure for the size of the polynomial should not be the degree, but the leading monomial. In this setting the footprint bound [25] becomes a natural enhancement of the factor theorem providing a bound on the number of common roots of general multivariate polynomials which is sharp for all finite Cartesian product point-sets. As our main contribution, by using methods from the theory of error-correcting codes we establish a natural formulation of the interpolation theorem to the case of common roots of multivariate polynomials. In short the two theorems reduce to the same result, but for dual spaces, establishing the unification requested in [44]. We leave it for further research to possibly establish similar interpolation results taking one or more of the various concepts of multiplicity of multivariate polynomials into account.
Keywords: Affine roots, footprint bound, generalized Alon-Füredi bound, interpolation, multivariate polynomials, polynomial method
@@ 1 Introduction
For a univariate non-zero polynomial over an arbitrary field the number of roots is at most equal to its degree, and if the field is algebraically closed and if roots are counted with multiplicity then equality holds. This result is in close correspondence with the fact that the interpolation problem of finding a polynomial of lowest possible degree evaluating to zero on a finite point-set boils down to calculating the product of linear terms each having one of the desired elements as root, and by repeating terms one can also deal with multiplicity. The situation is very much in contrast to that of multivariate polynomials. Firstly, a multivariate polynomial ring is not a principal ideal domain, and therefore the considered interpolation problem can not in general be about multiplying linear factors. Secondly, a non-constant multivariate polynomial often has infinitely many roots and always does if the field is algebraically closed, and therefore if the aim is to consider a case of finitely many roots then one either needs to restrict to roots from a prescribed finite point-set or one must study mutual roots of a set of polynomials which together define a zero-dimensional ideal.
@@
    Given some measure on the size of polynomials consider the class of all those of size within some given interval. What is the largest number $a$ such that for any point-set $A$ of cardinality equal to $a$ there exists $k$ linearly independent polynomials within the class evaluating simultaneously to zero on $A$, where $k$ is some adequate integer?
One often used measure on the size of a polynomial is its total degree in which case one may for instance in (I) and (II) consider the class of all polynomials of a given prescribed total degree. A much more fine grained measure, however, would be that of the leading monomial according to some fixed monomial ordering. The contribution of the present paper is to demonstrate that by applying the latter mentioned measure one is able to obtain for multivariate polynomials answers to (l) and (II) which are as strongly related to each other as are the answers in the univariate case. Recall that for univariate polynomials there is only one choice of monomial ordering and the leading monomial of a polynomial $F(X)$ of degree $d\geq 0$ equals $X^{d}$, hence the two concepts are basically the same. In contrast, for multivariate polynomials there are infinitely many monomial orderings and for many of these the total degree of the leading monomial does not even need to be equal to the total degree of the polynomial under consideration. Considering leading monomial(s) is not only as natural a choice for our study as considering total degree(s); as shall be demonstrated it is the key to achieve the above mentioned enhancement and connection. Even more importantly, using the leading monomial rather than the total degree provides much more fine-grained information which is sharp in several cases.
In his exposition [44] Tao treats problems (I) and (II) from the above paragraph applying the total degree as the measure of the size of polynomials and mostly concentrating on single polynomials rather than arbitrary sets of polynomials. He discusses on the one side Alon’s famous combinatorial Nullstellensatz [1, 33], the Alon-Füredi bound [2], the Schwartz-Zippel bound [47, 41] (as well as Dvir’s generalization taking multiplicity into account [14, 15, 16]) and on the other side the below standard result [44][Lem. 1] on interpolation; with the aim of solving problems in extremal combinatorics:
@@
Let $\mathbb{F}$} be a field, let $m\geq 1$ be an integer, and $d\geq 0$. If $A\subseteq{\mathbb{F}}^{m}$ has cardinality less than ${d+m}\choose{m}$, then there exists a non-zero polynomial $F\in{\mathbb{F}}[X_{1},\ldots,X_{n}]$ of degree at most $d$ having $A$ as roots.
@@
The theory of commutative algebra and algebraic geometry is very rich and indeed Tao in addition to the above results treats a cornucopia of deep results such as Bezout’s Theorem, Stepanov’s Method, the Hasse Bound and many more [4, 8, 30, 45, 34, 42, 46], as well as provide interpolation counterparts in more cases.
@@
Our contribution is to address the problem of common roots of general polynomials where the only assumption is that no two of them have the same leading monomial which is no restriction as we can always apply the division algorithm [10][Sec, 2, Par. 3]] without effecting the set of common roots. Similarly, we consider any finite point-sets without any restriction on their structure. Therefore, our answer to (II) should be compared to Proposition 1 and our answer to (I) should be compared to the Schwartz-Zippel bound, the Alon-Füredi bound and the generalized Alon-Füredi bound [7]. In both cases the material of the present paper can be viewed as improvements and of course also as enhancements as we in contrast to the mentioned bounds do not restrict to single polynomials but rather consider arbitrary finite sets of linearly independent polynomials.
@@
In the present paper we give refined answers to both (I) and (II) taking the leading monomial into account and we enhance the methods to treat the case of common roots of a set of more polynomials. Knowing the leading monomial corresponds to having important information regarding the most crucial coefficient and thereby our insight addresses the remark by Tao. More importantly, we demonstrate a very strong relationship between our answers to (I) and (II) which is then our proposal for the requested useful unification, except we do not treat multiplicity.
Our answer to (I) is basically to recall some instances of the footprint bound [11, 31, 25], but we choose to present in this paper as an answer to (I) also an enhanced version of the Feng-Rao bound for primary codes [3, 22, 27] demonstrating that this bound works over arbitrary fields rather than only over finite fields. The latter framework is more general as our footprint bound supported answer to (I) can be derived as an easy corollary to the Feng-Rao bound for primary codes. More importantly, the natural framework for us to prove our answer to (II) is to combine some basic linear algebra results stated by Forney in [21] with an enhanced version of the Feng-Rao bound for dual codes [18, 19, 20, 32, 28] (although we could have also chosen to replace the latter with an enhanced version of the order bound [32] from order domain theory which would represent an intermediate level between the polynomial ring description and the pure linear algebra description). As shall become clear from our analysis, in a certain sense problems (I) and (II) are basically the same, but stated for dual spaces. This duality is also what allows us to prove that the answer to (II) is sharp in the finite field case. We stress that our description at linear code level should not be considered as a detour for us to treat the case of general sets of multivariate polynomials. In our view the description at linear code level is the most fundamental description with the potential of supporting future investigations beyond the level of general sets of multivariate polynomials and general affine point-sets.
@@
The paper is organized as follows. We start in Section 2 by presenting in Theorem 5 our main finding regarding (II) and by recalling various versions of the footprint bound which address (I). In that section we also in Remark 11 provide a description of the relationship between these results. To prepare for the proof of the first mentioned result we then in Section 3 treat related problems in the language of error-correcting codes over arbitrary fields. In Section 4 as a corollary we then obtain the proof of our answer to (II) which also demonstrates the strong relationship between (I) and (II), and which implies that our answer to (II) is sharp in the case of finite fields. Section 5 contains concluding remarks.
@@ 2 Main results on multivariate polynomials
Our main finding is the following enhancement of Proposition 1 which surprisingly is the best possible general result as for finite fields we can show it to be sharp. Observe that if $\mathbb{F}$} is a finite field with $q$ elements, in which case we write ${\mathbb{F}}={\mathbb{F}}_{q}$, and if $q\leq d$ then Proposition 1 does not give us any real information as all points in ${\mathbb{F}}^{n}$ are then trivially roots of each of the polynomials $X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}$. The below theorem is prepared to also give non-trivial insight in such cases.
@@
For multivariate monomials $N_{1},\ldots,N_{s}$ write
@@
constitutes a basis for ${\mathbb{F}}_{q}^{n}$ as a vector-space over ${\mathbb{F}}_{q}$. In particular for any polynomial $G$ with support not being a subset of $\Box$ there exists a polynomial $H$ with support being a subset of $\Box$ in such a way that $G$ and $H$ evaluates similarly at ${\mathbb{F}}_{q}^{m}$ and such that if $H$ is not the zero-polynomial then ${\mbox{lm}}(H)\preceq{\mbox{lm}}(G)$ holds for any choice of monomial ordering $\prec$. Here, ${\mbox{lm}}()$ denotes the leading monomial. More concretely, $H$ can be obtained from $G$ by dividing the latter with $\{X_{1}^{q}-X_{1},\ldots,X_{m}^{q}-X_{m}\}$. Therefore it is no restricting to consider in the finite field case only polynomials with support in $\Box$.
Given a polynomial $F$ in the following by ${\mbox{Supp}}(F)$ we shall denote the support of it, i.e. ${\mbox{Supp}}(F)$ is the set of monomials that appear with non-zero exponents in the description of $F$. We note that later in the paper we shall also use the notion ${\mbox{Supp}}()$ in another meaning, when we apply it to vectors and vector-spaces.
@@
We note that the right hand side of (2) equals the dimension of the q-ary Reed-Muller code ${\mathcal{RM}}_{q}(d,m)$ [13, Lem. 2.4.6]).
We next address (I) by recalling the footprint bound on the number of common roots of any given set of polynomials over arbitrary field [11, Ch. 2, Thm. 2.10] and [31, 25]. To the best of our knowledge this is the most general and sharp bound for arbitrary sets of polynomials over arbitrary point-sets. Let $\prec$ be a monomial ordering on the set of monomials in $m$ variables, and consider an ideal $I\subseteq{\mathbb{F}}[X_{1},\ldots,X_{m}]$. The footprint of $I$ with respect to $\prec$ is
@@
The crucial observation is that $\{M+I\mid M\in\Delta_{\prec}(I)\}$ constitutes a basis for ${\mathbb{F}}[X_{1},\ldots,X_{m}]/I$ as a vector-space over $\mathbb{F}$}, [10, Ch. 5, Sec. 3, Prop. 4], which is seen by applying the division algorithm.
@@
Assume first that $V_{\mathbb{F}}(I)$ is finite and write $V_{\mathbb{F}}(I)=\{\beta_{1},\ldots,\beta_{n}\}$. Consider the evaluation map ${\mbox{ev}}:{\mathbb{F}}[X_{1},\ldots,X_{m}]/I\rightarrow{\mathbb{F}}^{n}$ given by ${\mbox{ev}}(F+I)=(F(\beta_{1}),\ldots,F(\beta_{n}))$. By interpolation arguments this map is surjective and the first part of the theorem follows. To see the last part we use the fact [10][Ch. 4, Sec. 2, Thm. 7] that ${\mathcal{I}}(V_{\mathbb{F}}(I))=I$ whenever $I$ is radical and $\mathbb{F}$} is algebraically closed. Aiming for a contradiction we finally assume that $V_{\mathbb{F}}(I)$ is infinite and choose $\{\beta_{1},\ldots,\beta_{n}\}\subseteq V_{\mathbb{F}}(I)$ with $n>\#\Delta_{\prec}(I)$. But then ev cannot be surjective. ∎
@@
$I=\left\langle G_{1}=\prod_{b\in A_{1}}(X_{1}-b),\ldots,G_{m}=\prod_{b\in A_{m}}(X_{m}-b)\right\rangle$ (3)
@@
$\Delta_{\prec}(I)=\{M{\mbox{ is a monomial }}\mid\deg_{X_{1}}M<\#A_{1},\ldots,\deg_{X_{m}}M<\#A_{m}\}.$ (4)
Note, that if $A_{1}=\cdots=A_{m}={\mathbb{F}}_{q}$, then this set equals $\Box$ where the latter is described in Definition 3.
@@
Let $A_{1}\times\cdots\times A_{m}$ be a finite point-set and consider an arbitrary monomial ordering $\prec$. Let $I$ and $\Delta_{\prec}(I)$, respectively, be as in (3) and (4), respectively. Consider pairwise different monomials $M_{1},\ldots,M_{t}\in\Delta_{\prec}(I)$. For general set of $k\leq t$ polynomials $F_{1},\ldots,F_{k}\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ with ${\mbox{Supp}}(F_{i})\subseteq\Delta_{\prec}(I)$, $i=1,\ldots,k$ and leading monomials being pairwise different all belonging to $\{M_{1},\ldots,M_{t}\}$ the number of non-common roots of $F_{1},\ldots,F_{k}$ in $A_{1}\times\cdots\times A_{m}$ is at least
@@
where $N:=X_{1}^{\#A_{1}-1}\cdots X_{m}^{\#A_{m}-1}$. This bound is sharp.
@@
We only need to demonstrate sharpness. To this end write $a_{i}=\#A_{i}$, $A_{i}=\{b_{1}^{(i)},\ldots,b_{a_{i}}^{(i)}\}$ for $i=1,\ldots,m$. To each monomial $N\in\Delta_{\prec}(I)$ we associate the polynomial
@@
which clearly has leading monomial $N$. Then for any set of monomials $\{N_{1},\ldots,N_{s}\}\subseteq\Delta_{\prec}(I)$ the number of non-common roots of $H_{N_{1}},\ldots,H_{N_{s}}$ is precisely $\sigma(N_{1},\ldots,N_{s})$. ∎
A special case of Corollary 10 is when $A_{1}=\cdots=\#A_{m}=:a$ and there is only one polynomial $F$, the only information of which we have is its total degree $\deg F=d$. Writing $d=(\sum_{i=1}^{v}(a-1))+\ell$ where $0\leq\ell<a-1$ we obtain the following upper bound on the number of roots
@@
which is an incidence of the Alon-Füredi bound [2] that in its most general version allows for $A_{1},\ldots,A_{m}$ to be of different sizes. The general version of the Alon-Füredi bound can be deduced from Corollary 10 in a similar manner.
@@
Let ${\mathbb{F}}={\mathbb{F}}_{8}$, i.e. the field with $8$ elements, and consider $I=\langle X_{1}^{8}-X_{1},X_{2}^{8}-X_{2}\rangle\subseteq{\mathbb{F}}_{8}[X_{1},X_{2}]$. Clearly, $V_{\mathbb{F}_{8}}(I)={\mathbb{F}}_{8}^{2}$. We have $\Delta_{\prec}(I)=\{X_{1}^{i_{1}}X_{2}^{i_{2}}\mid 0\leq i_{1},i_{2}<8\}$ regardless of the choice of monomial ordering $\prec$. This set of course is the set $\Box$ from Definition 3. Assume $M_{1}=X_{1}^{3}X_{2}$, $M_{2}=X_{1}^{2}X_{2}^{2}$, and $M_{3}=X_{1}X_{2}^{3}$ and that $\prec$ is the degree lexicographic ordering with $X_{1}\prec X_{2}$ which implies that indeed $M_{1},M_{2},M_{3}$ are strictly consecutive. We have $\mu(M_{1})=\mu(M_{3})=8$, $\mu(M_{2})=9$, $\mu(M_{1},M_{2})=\mu(M_{2},M_{3})=11$, $\mu(M_{1},M_{3})=12$, and $\mu(M_{1},M_{2},M_{3})=13$. From Theorem 5 we conclude that for each $A\subseteq{\mathbb{F}}_{8}^{2}$ of size at most $7$ there exist polynomials $F_{1},F_{2},F_{3}$, ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, ${\mbox{lm}}(F_{3})=M_{3}$ having all elements of $A$ as common roots. We similarly conclude that whenever $A$ is of size at most $10$ then there exist $F_{1},F_{2}$ with ${\mbox{lm}}(F_{1})\neq{\mbox{lm}}(F_{2})$, ${\mbox{lm}}(F_{1}),{\mbox{lm}}(F_{2})\in\{M_{1},M_{2},M_{3}\}$ with all elements of $A$ as common roots. Finally for any $A$ with #A<13\#A<13 there exists $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{2},M_{3}\}$ such that $F(a)=0$ for all $a\in A$. In the other direction $\sigma(M_{1})=\sigma(M_{3})=35$, $\sigma(M_{2})=36$, $\sigma(M_{1},M_{2})=\sigma(M_{2},M_{3})=41$, $\sigma(M_{1},M_{3})=45$ and $\sigma(M_{1},M_{2},M_{3})=46$ which by Corollary 10 tell us that any polynomial $F$ with ${\mbox{lm}}(F)\in\{M_{1},M_{3}\}$ has at most 6435=2964-35=29 roots and that if ${\mbox{lm}}(F)=M_{2}$ then it has at most 6436=2864-36=28 roots. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{2}$ or if ${\mbox{lm}}(F_{1})=M_{2}$ and ${\mbox{lm}}(F_{2})=M_{3}$ then $F_{1}$ and $F_{2}$ have at most 6441=2364-41=23. If ${\mbox{lm}}(F_{1})=M_{1}$ and ${\mbox{lm}}(F_{2})=M_{3}$ the number of common roots is at most 6445=1964-45=19. Finally $F_{1},F_{2},F_{3}$ with ${\mbox{lm}}(F_{1})=M_{1}$, ${\mbox{lm}}(F_{2})=M_{2}$, and ${\mbox{lm}}(F_{3})=M_{3}$ have at most 6446=1864-46=18 roots in common.
The similarity in expression (1) and expression (5) suggests that there is a strong relationship between Theorem 5 and Corollary 10 which we now explain. Note, however, that there in contrast to Theorem 5 is in Corollary 10 no requirement that the monomials $M_{1},\ldots,M_{s}$ should be strictly consecutive.